a pair of fair dice is tossed. the probalility of getting at least a total of 4 is:
Answer by ewatrrr(24785)

36 total outcomes
P(x ≥ 4) = 33/36 = 11/12
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12